CURRICULUM VITAE - Oliver Gerler

Personal Details:

Full name:  Oliver Mikis GERLER
Date of birth: 17/09/73
Email:  oge@rockus.at
Address:  Schumanngasse 13

A-8010 Graz Austria

Phone number: +43-664-9204150
Nationality:  Austrian

Education:

1997-2003 University of Graz, Austria

Diplomstudium der Astronomie

qualification obtained at completing 1^{st} part of degree in April 2003:

1. Diplomprüfung aus Astronomie (comparable to Bachelor of Science in Astronomy)

2002 International Space University, Pomona, California, USA

Summer Session Program 2002

The curriculum covers the principal space related fields, both non-technical and technical and ranges from policy and law, business and management and humanities to life sciences, engineering, physical sciences and space applications.

1993-2001 University of Technology, Graz, Austria

Diplomstudium der Elektrotechnik, Studienzweig Elektronik und Nachrichtentechnik

qualification obtained at completing degree in September 2001: Diplomingenieur aus Elektrotechnik (Master of Science in Electrical Engineering, electronics and telecommunications) My Master’s thesis describes a process of reloading tasks on embedded DSP systems at runtime under special considerations of low overhead.

The academic year 1999/2000 was spent at the University of Limerick, Ireland, as an Erasmus student with successfully passed exams on Digital Signal Processing, Real Time Systems, Optical Fibre Communications, Expert Systems and Artificial Intelligence.

June 1992 Matura (qualification at graduation of secondary education):

German (4), English (3), Mathematics (1), Physics (1), Descriptive Geometry (4) (1 very good, 2 good, 3 reasonable, 4 sufficient, 5 failed)

1984-1992 Secondary (Academic) School

BG/BRG Lichtenfelsgasse 3-5, Graz, Austria

1980-1984 Primary School

Volksschule Waltendorf, Graz, Austria Relevant Employment:

since 11/2012 Systems Architect, Product Owner, Maxim Integrated GmbH, Lebring/Graz, Austria Systems Architecture and Simulations of automotive radio transceivers. Requirements Elicitation and Management with in-depth customer relations.

01/2011-10/2012 Systems Design Engineer, NXP GmbH, Gratkorn, Austria Systems design of digital receiver building blocks.

10/2009-12/2010 Project Manager, Systems Engineer, SensorDynamics AG, Lang/Lebring, Austria Systems design of mixed-signal LF transmitters including sensor interfaces. Management of customer developments within tight scheduling and limited resources. Customer relations interface. Management of ASIC development procedures and tasks ranging from packaging to lab evaluation and production test.

05/2006-09/2009 Digital Design Engineer, SensorDynamics AG, Lang/Lebring, Austria Simulation/Implementation of power- and area-optimised digital signal-processing units for RF transceivers. Transfer of designs from systems to implementation level. Design and implementation supervising of functional verification system for laboratory and production ASIC tests.

01/2004-04/2006 Software Systems Design & Quality Management, Magna Powertrain, Lannach, Austria Team Leader Software Development. Design of embedded operating systems for automotive applications. Definition and integration of structured software systems design and development processes and according quality management system. Management of software development including co-ordination of clients and suppliers. Design of network and server infrastructure for software requirements management (ISO 15504, SPiCE).

03/2004-07/2004 Lecturer, Fachhochschule Kapfenberg, Kapfenberg, Austria (parallel to main job) ”Networking 3: Application protocols and heterogeneous systems“

12/2001-11/2003 Hardware Developer, Austrian Academy of Sciences, Space Research Institute, Graz, Austria Design and development of electron gun to be flown on NASA mission. Evaluation of FPGA soft-cores for future use as instrument controller.

Positions of Responsibility and Interests:

Editor, writer and researcher for Final Report of the Astrobiology Design Project produced during International Space University Summer Session Programme 2002 in Pomona, California.

Successful application of my Concurrent Editing design concept at above session, simplifying to organise writers/researchers under time constraints without restricting their choice of writing environment. Extended usage of cross-platform script-utilities, mostly written by myself, allows to automate the document compile process. The concept comfortably handled a number of writers and was able to help achieve timely completion of the project.

Interest in wildlife conservation and environment protection on the technological level.

Interest in designing analogue and digital circuits and the corresponding software for private embedded systems. Following development on telecommunications market.

Initiator and administrator of the International Space University alumni mailing list, serving as a central point of communication to the numerous people subscribed. I am also involved in arranging alumni reunions which are mainly organised by use of that list.

Editor and writer for the first Amiga online magazine amigascene.at.

Interest in literature, fantasy, science fiction and technological as well as sociological literature on the open source movement.

Skills:

Language Proficiency: Fluent in German and English

Transferable Skills: I need to see the big picture. Understanding the context of applications is crucial for me. As a result, I live the systems approach.

In the course of my studies and employment I acquired skills such as presentation, organisation and group management, communication, independence and research.

The application of the Concurrent Editing concept during the International Space University Summer Session 2002 has taught me to work well in a team, both in leading and support roles. This also required me to act as an interface between the 50 writers and the system to meet deadlines.

Electronics: Simulation of mixed signal automotive transceiver projects together with transfer of systems design to bit-true simulation and test-bench data generation

Implementation and verification of digital signal-processing units in VHDL, Verilog and SystemC

Development, implementation and review of both hardware and software for electronic control unit designs in the automotive market

Design and prototyping of analogue and digital electronics for private projects

IT: Daily usage of Matlab/Simulink for Simulation and Cadence design tools for implementation and test-benches

Very good knowledge of operating systems such as AmigaOS, Linux, SGI IRIX, NeXTstep, netBSD, SunOS, AIX, Windows, together with in-depth knowledge of networks (installation, administration and applications)

Several years of experience of operation of personal network server farm; applying firewalls, traffic shaping, security implementation, mail server, web server, database server; desktop machine setup

Detailed knowledge of hardware specifics of Amiga, NeXT, PC

Knowledge of automotive development tools such as CANape, CANoe, CANalyzer

Experience in C, Assembler (m68k, TMS320C40, TMS320C6000), Perl, PHP, HTML, CSS

Knowledge of L^{A}TEX type setting environment

Extended knowledge of Windows and Office in several versions Scholarships, Grants and Projects:

09/2015 griehser.at Sicherheit ganz oben Graz, Austria Höhenfacharbeiter, Level 1 ISO 22846

2013/2014 Maxim Integrated Products San Jose, California, USA Maxim university “Requirements Management and Elicitation”

2013/2014 Maxim Integrated Products San Jose, California, USA Maxim university “Mixed-Signal System-On-A-Chip Development”

04/2005 HPS Wien, Austria advanced training “Effective Business Presentations”

01/2005 Magna Powertrain Lannach, Austria advanced training “Führen für Meister und Gruppenleiter” (team leading)

08/2004 Swiss Association for Quality Basel, Switzerland advanced training “Certified Software Tester, Foundation Level - ASQF-Certified-Tester”

07/2002-08/2002 International Space University Summer Session Programme Pomona, California, USA space related hands-on experience in: system analysis and design, satellite applications, life sciences in space, business and management, financing, international law and anthropology.

07/2001 European Space Agency Summer School Alpbach 2001 Alpbach, Austria “Satellite Navigation Systems for Science and Application”

03/2001 TMS320C6000 Workshop Graz, Austria Texas Instruments Workshop on programming high-end DSPs.

09/1999-05/2000 Erasmus Student Exchange University of Limerick, Ireland

focus on signal processing

03/1998-02/1999 XilinX CPU Institute for Informatics University of Technology Graz, Austria Development of a CPU core based on XilinX FPGAs with hardware multitasking and support for specialised solutions.

01/1998-07/1999 amigascene.at online editor, writer and hardware tester for the first Austrian Amiga online magazine.

03/1993 Strahlenschutzleistungsbewerb Seibersdorf, Austria

Course on anti-radiation precautions.

01/1992-06/1992 11. Österreichische Physik-Olympiade Spittal/Drau, Austria (Austrian Physics Contest)

Successful completion of Styrian and Austrian contests and participation in preparation weeks.